Use the given degree of confidence an sample data to find a confidence interval for the population standard deviation σ. Assume that the population has a normal distribution.

List the critical values

Number of ounces of coffee in a 13.7oz bottle of Strabucks Mocha drink (measured cold at 40 degrees F): 95% confidence; n = 18, x̅ = 13.8oz, s = 0.03oz
Answer #1

Solution: Given that 0=18, X=1318, s=0.03 At 95%. confidence Level 2=1-95% -1-0.95 -0.05 M = 0.05/2 = 0.025 df = n-12 18-1 =22 (18-130.034 (18-130.03 <a < j 30:191 4.564 0.0153 41. 0.0153 <a< -- 30:191 7564 -Jo.co 0.0005064 <a j 0.002023 0.0225<a <0